What is the general rule for the following terms{2,1,4,4,6,9,8}
Answer by Edwin McCravy(20056)

a1 = 1+1 = 2
a2 = 2²/4 = 1
a3 = 3+1 = 4
a4 = 4²/4 = 4
a5 = 5+1 = 6
a6 = 6²/4 =9
a7 = 7+1 = 8

an = [n mod 2](n+1) + [(n+1) mod 2](n²/4) 

Edwin